# Asymptotics of the d’Arcais Numbers at Small $k$

### v2

@@ 2 Main results
In [14], Hong and Zhang announced their proof of Heim and Neuhauser’s conjecture for all $k$, provided $n$ is sufficiently large. They have explicit bounds on $n$, which we do not provide. Presumably, their proof does not apply to $k=2$ . (For example in their Lemma 2.3, they stipulate $n\geq k^{4}\log^{4}k$ which is undefined if $k=1$.) The first part of our Corollary gives a sequence of counterexamples: large powers of primorials $n=(p_{m}\#)^{r}$ for a sequence $m$ and $r$ approaching $\infty$. (See Appendix B.) In terms of the other values $k=3,4,\dots$, the added value of our result is the precise limit inferior.
@@ Appendix B A review of the divisor functions
If we take a sequence of increasingly high powers of increasingly large primorials, then we can approach the limits where both inequalities are saturated. That is how we obtain the limit inferior result.

### v3

@@ 2 Main results
In [14], Hong and Zhang announced their proof of Heim and Neuhauser’s conjecture for all $k$, provided $n$ is sufficiently large. They have explicit bounds on sufficient sizes for $n$, for their bounds to apply. That is valuable and we do not provide that. It seems to us that their proof does not apply to $k=2$ . (For example in their Lemma 2.3, they stipulate $n\geq k^{4}\log^{4}k$ which is undefined if $k=2$.)
Remark 2.6
The most interesting regime is $k=xn^{1/2}$ for $x\in(0,1)$. This is the typicality regime, as Abdesselam points out in [1]. In that regime, he obtained a full asymptotic expansion such that $A(\ell,n,k)\sim\widetilde{A}(\ell,n,k)$ as $n\to\infty$, for an explicit formula found by himself. He can also handle $\ell=3,4,\dots$ as well as $\ell=2$. In a separate paper, the author considered just $\ell=2$ (as in this note) and considered large deviations assuming $1\ll k$ and $k/n<1-\epsilon$ for a fixed $\epsilon>0$ [24]. The interesting regime of $k/n\to 1$ is also treated by Raghavendra Tripathi [25] where he proves uniform log-concavity in an interesting way.
@@ 5 Outlook and open problems
Acknowledgments
I am grateful to Raghav Tripathi for an interesting discussion with useful comments. I am grateful to Malek Abdesselam for introducing me to this topic.
@@ Appendix B A review of the divisor functions
Remark B.1
If we take a sequence of increasingly high powers of increasingly large primorials, then we can approach the limits where both inequalities are saturated. That is how we obtain the limit inferior result. The limit inferior claim of Corollary 2.4 comes from taking a sequence of increasingly high powers of increasingly large primorials. Then we can approach the limits where both inequalities are saturated. That is how we obtain the limit inferior result. Consider $n=(p_{m}\#)^{r}$ for a sequence $m$ and $r$ approaching $\infty$.
Remark B.2
The most surprising result is that $\big(\alpha(n,2)\big)^{2}/\big(\alpha(n,1)\alpha(n,3)\big)$ has limit inferior equal to 0. Note that this means $\big(A(2,n,2)\big)^{2}/\big(A(2,n,1)A(2,n,3)\big)$ has limit inferior equal to 0 as well, since the latter merely equals the former times $(3!1!)/(2!)^{2}$. Similarly, for $k=3,4,\dots$ since the limit inferior of $\big(\alpha(n,k)\big)^{2}/\big(\alpha(n,k-1)\alpha(n,k+1)\big)$ is greater than 1, so is the limit inferior of $\big(A(2,n,k)\big)^{2}/\big(A(2,n,k-1)A(2,n,k+1)\big)$ since $((k+1)!(k-1)!)/(k!)^{2}=(k+1)/k>1$.